OpenStudy (anonymous):

What is the equation of the line that passes through (–2, –3) and is perpendicular to 2x – 3y = 6?

OpenStudy (austinl):

Okay, are you familiar with slope-intercept form?

OpenStudy (anonymous):

I am.

OpenStudy (anonymous):

first you need to find the slope of the given equation. Can you do that ?

OpenStudy (arabpride):

u have to plug in the x coordinate into where it says 2x, then y coordinate to where it says -3y......nd work with that

OpenStudy (austinl):

Okay, I would first put it in that form. You essentially just solve for y.

OpenStudy (austinl):

And in doing so, you find the slope as well. The term that ends up in front of x.

OpenStudy (anonymous):

Not sure if I did this right, Y=(2/3)x-2?

OpenStudy (austinl):

\(2x-3y=6\) \(-3y=-2x+6\) \(\displaystyle y=\frac{2}{3}x-2\) Good job, now what is the slope?

OpenStudy (anonymous):

(2/3)

OpenStudy (austinl):

Correct. Now for a line to be perpendicular to another, it's slope is the negative reciprocal. Do you know how to do that?

OpenStudy (anonymous):

-(3/2) right?

OpenStudy (austinl):

Yes! So now we have this basic equation. \(\displaystyle y=-\frac{3}{2}x+b\) We need to solve for b, do you know how we would do this?

OpenStudy (austinl):

I mean, figure out what b is. There is a difference :P

OpenStudy (anonymous):

I think I know the answer \[Y=-\frac{ 3 }{ 2 } x-6?\]

OpenStudy (austinl):

That is correct! Good job!